### v1

# Sample Compression for Continual Learning

Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees for the continual learning setup. In this paper, we present a new method called ‘Continual Pick-to-Learn’ (CoP2L), which is able to retain the most representative samples for each task in an efficient way. The algorithm is adapted from the Pick-to-Learn algorithm, rooted in the sample compression theory. This allows us to provide high-confidence upper bounds on the generalization loss of the learned predictors, numerically computable after every update of the learned model. We also empirically show on several standard continual learning benchmarks that our algorithm is able to outperform standard experience replay, significantly mitigating catastrophic forgetting.
Machine Learning, ICML
1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999; Goodfellow et al., 2014). In order to cope with forgetting, various methodologies have been developed, such as the regularization based approaches (e.g., Kirkpatrick et al., 2017), architectural approaches (e.g., Rusu et al., 2016), or rehearsal based approaches (e.g., Rolnick et al., 2019).
It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019). However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. (2024). P2L is a meta algorithm that selects a compression set, which amounts to express the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory (Floyd & Warmuth, 1995; Campi & Garatti, 2023; Laviolette et al., 2005) in mind. By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor.
In this paper, we propose ‘Continual Pick-to-Learn‘ (CoP2L), an algorithm that leverages the sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly, and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contribution overall can be summarized as follows :
  • We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
  • CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks.
  • We experimentally show that CoP2L significantly mitigates forgetting, while also being able to obtain better performance compared to vanilla replay, in the settings where the replay buffer is small.
1.1 Related Work
Practical Continual learning
In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories which include regularization approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018), architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya & Lazebnik, 2018; Wang et al., 2022) and rehearsal based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017; Buzzega et al., 2020). In this paper, we use experience replay as a baseline as it is shown in the literature to be a very strong continual learning baseline to mitigate forgetting (Libera et al., 2023).
Theoretical Continual Learning
There have been several works in theoretical continual learning in the recent years. For example, Bennani & Sugiyama (2020); Doan et al. (2021) study the generalization error obtained with Neural Tangent Kernel (NTK) models when the model is trained with Orthogonal Gradient Descent (Farajtabar et al., 2020).
Yin et al. (2020) provides generalization bounds for regularization-based continual learning methods. Asanuma et al. (2021) studies the impact of task similarity in the student-teacher setup. Li et al. (2022) provides a sample complexity bound in the context of continual representation learning for an adapted PackNet (Mallya & Lazebnik, 2018). In Chen et al. (2022), the authors provide a theoretical bound on memory using PAC learning theory. These works however do not provide bounds that are computable in practice.
A line of work also investigates the generalization of continual learning models when linear models are employed. In Goldfarb & Hand (2023), the authors investigate the effect of overparametrized linear models. In Evron et al. (2022); Lin et al. (2023), the authors analyze the generalization behavior of linear regression models, even though Lin et al. (2023) conjectures that their results could be extended to neural networks as well.
Different from all the listed works above, in this paper, we provide a practical non-trivial generalization bound that is computed on the training set, and predicts the test error for each task separately. Furthermore, also unlike the approaches above, our bound is applicable to any neural network architecture and does not have limitations on the continual learning setup that is employed.
Sample Compression
The sample compress theory has been shown effective for providing tight generalization bounds (Laviolette et al., 2005; Marchand et al., 2003; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Shah, 2007). However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks (Campi & Garatti, 2023; Paccagnan et al., 2024). In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
2 Background
2.1 Continual Learning
Notation.
Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$.
We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$.
Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}(\theta_{T})\ \coloneqq\operatorname*{\mathbf{E}}_{({\mathbf{x}},y)\sim{\mathcal{D}}_{t}}\ell(\theta_{T},{\mathbf{x}},y)\,.$ (1)
The empirical loss counterparts on observed training samples are given by
^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999). Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
${\mathcal{F}}_{t}^{\ \text{replay}}(\theta_{t})\ \coloneqq\ \widehat{\mathcal{L}}_{S_{t}}(\theta_{t})+\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}\ell(\theta_{T},{\mathbf{x}}_{j},y_{j})\,,$ (3)
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as the baseline as it’s shown to be a simple and very effective continual learning strategy (Libera et al., 2023; Mai et al., 2020).
2.2 Sample Compression Theory
Given a dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $A(S)$ on the condition that the predictor can be provably represented as a function of a small subset of the dataset, called the compression set, and an additional source of information, called the message. If this is the case, we call $A(S)$ a sample-compressed predictor.
The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that we have
$\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
The compression set is defined such that
$S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ (5)
We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S_{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor. Although generally defined as a binary sequence, the message set can be defined by any set of countable sequences of symbols. Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 00 to $\infty$, constructed using the alphabet $\Sigma$. For all $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
Given a learning algorithm $A$ and a dataset $S$, to prove that the predictor $A(S)$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\bigcup_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message that are used to represent $A(S)$. Given a dataset $S$, we have $\mathscr{C}(S)=(S^{\operatorname{\mathbf{i}}},\mu)$. Then, the reconstruction function $\operatorname{\mathscr{R}}:\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. Both the reconstruction function and the compression function must be deterministic and data-independent.
The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following the work of Marchand & Sokolova (2005), we consider $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{2}}(k+1)^{-2}$. Given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
We now present the sample compression bound of Bazinet et al. (2025). We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that don’t belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\mathbf{i}\in\operatorname{\mathscr{P}}(n),\,\mu\in\mathscr{M}(\operatorname{\mathbf{i}}):$
$\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)))\leq\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and
$\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})].$
Given the parameters $\theta$ of a reconstructed predictor, i.e., $\theta{=}\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, Theorem 2.1 states that the $\mathrm{kl}$-discrepancy between the true risk ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\theta)$ and the empirical risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ is upper bounded by the complexity term $\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$. From this result, one can compute an upper bound on the true risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ by inverting the Kullback-Leibler divergence. Indeed, Theorem 2.1 can be rewritten as
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta)\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta),\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta))$ (6)
with
$\mathrm{kl}^{-1}(q,\varepsilon)\ =\ \operatorname*{arg\,sup}_{0\leq p\leq 1}\{\mathrm{kl}(q,p)\leq\varepsilon\}.$ (7)
On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase. On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
2.3 Pick-to-Learn
To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) (Marchand & Shawe-Taylor, 2002; Marchand et al., 2003; Marchand & Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al. (2024) as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm modifies the training loop of the predictor, by iteratively choosing datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of Pick-To-Learn in Algorithm 1.
Algorithm 1 Original Pick-To-Learn (P2L)
0:  $S=\{({\mathbf{x}}_{i},y_{i})\}_{i=1}^{n}$Training set}
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
0:  $\gamma$Stopping criteria}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S}\ell(\theta_{0},{\mathbf{x}},y)$
3:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\geq\gamma$ do
4:     $k\leftarrow k+1$
5:     $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y})\}$
6:     $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
7:     $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)$
8:  end while
9:  return $\theta_{k},C_{k}$.
10:  {Learned parameters and compression set}
3 Sample Compression for Continual Learning
3.1 A New Training Scheme
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer, thus mitigating the forgetting by adding a well-chosen datapoint.
Algorithm 2 Modified Pick-To-Learn (mP2L)
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
0:  $S=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{n}$Training set (with weights)}
0:  $B^{\star}=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{m}$Buffer set (with weights)}
0:  $\gamma$Stopping criteria}
0:  $K^{\star}$Maximum number of iterations}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $S^{\star}\leftarrow S\cup B^{\star}$
3:  $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}}\ell(\theta_{0},{\mathbf{x}},y)\cdot w$
4:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\cdot\overline{w}\geq\gamma$ and $k\leq K^{\star}$ do
5:     $k\leftarrow k+1$
6:     $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y},\overline{w})\}$
7:     $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
8:     $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)\cdot w$
9:  end while
10:  if $k<K^{\star}$ then
11:     $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$
12:  end if
13:  return $\theta_{k},C_{k}$.
14:  {Learned parameters and compression set}
Algorithm 3 Continual Pick-To-Learn (CoP2L)
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
1:  $B_{i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
7:      $B_{t}\leftarrow\mbox{sample}(\hat{S}_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
8:     $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
9:  end for
10:  return $\theta_{T}$. {Learned parameters}
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm CoP2L calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum size of the buffer. We present CoP2L in Algorithm 3.
We modified Pick-To-Learn in two significant ways. First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer.
The second modification leading to mP2L is the choice of the returned predictor. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior. Instead, to perform early stopping, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
$\Psi(S,\theta,C)$ $=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}\quantity(\theta),\ \tfrac{1}{|S\setminus C|}\log\quantity(\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta})),$ (8)
with $\mathrm{kl}^{-1}$ defined by Equation 7. Thanks to Theorem 2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with probability at least $1-\delta$.
Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L). Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$. It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks. The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) the bound function $\Psi(S,\theta,C)$ cannot be computed on the whole dataset, when only the compression set is given as input; (2) the sampling procedure sample$(S,m)$ cannot be reproduced in the reconstruction step. In the following section, we define the compression and reconstruction functions of CoP2L, which will give rise to sample compression bounds as presented in Theorem 3.1.
3.2 Compression and Reconstruction Scheme
To obtain generalization bounds for CoP2L, we need to prove that it can be compressed. To do so, we provide the compression and reconstruction functions of CoP2L. The compression function is presented in Algorithm 4 (provided in supplementary material) and the reconstruction function is presented in Algorithm 5.
Compression function.
In line with the prevalent literature, we presented the sample compression framework (Section 2.2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^1] The idea of using multiple compression sets appeared in Marchand & Shawe-Taylor (2002) and Marchand et al. (2003), but in a setting without a message Two challenges prevent CoP2L from being used as its own compression and reconstruction function as the original P2L.
The first challenge comes from the sampling of the buffer ${\mathcal{B}}$. When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message thus provide the task after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second one is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
After learning on $T$ tasks, the compression function provides the sample compression set $S^{\operatorname{\mathbf{i}}}$ and  $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\bigtimes\limits_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (9)
Reconstruction function.
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function and for CoP2L, the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is adressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
Algorithm 4 Compression function of CoP2L
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
1:  $B_{i}\leftarrow\emptyset,C_{i}\leftarrow\emptyset,C^{\mu}_{i}\leftarrow\emptyset,\mu_{1,i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset,\mu_{2}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:     $\mu_{2}^{t}\leftarrow|C^{\star}|$
7:     $C_{i}\leftarrow C_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
8:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
9:     $B_{t}\leftarrow\mbox{sample}(S_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
10:     for $i\in\{1,\ldots,t-1\}$ do
11:        for $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\in B^{\star}$ and $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\notin B_{i}$ do
12:           $C^{\mu}_{i}\leftarrow C^{\mu}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
13:           $C_{i}\leftarrow C_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
14:           $\mu_{1,i}^{j}\leftarrow\{t\}$
15:        end for
16:     end for
17:     $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
18:  end for
19:  return $\theta_{T},\{C_{t}\}_{t=1}^{T},\{C^{\mu}_{t}\}_{t=1}^{T},\{M_{t}\}_{t=1}^{T},N$. {Learned parameters, compression sets and messages sets}
Algorithm 5 Reconstruction function of CoP2L
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
0:  $C_{1},C_{2},\ldots,C_{T},\ C^{\mu}_{1},C^{\mu}_{2},\ldots,C^{\mu}_{T}$Compression sets}
0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T},\,\mu_{2}${Message sets}
1:  $B^{\star}\leftarrow\emptyset$
2:  for $t\in\{1,\ldots,T\}$ do
3:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in C_{t}}$
4:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\mu_{2}^{t})$
5:     $B_{t}\leftarrow(C_{t}\setminus C^{\star})\cup C^{\mu}_{t}$
6:     for $i\in\{1,\ldots,t-1\}$ do
7:        for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in C^{\mu}_{i}$ do
8:           if $\mu_{1,i}^{j}=t$ then
9:              $B_{i}\leftarrow B_{i}\setminus\{(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\}$
10:           end if
11:        end for
12:        $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
13:     end for
14:  end for
15:  return $\theta_{T}$. {Learned parameters}
3.3 Generalization Bound
The last subsection provided the compression and reconstruction functions of CoP2L. Thus, we know CoP2L provides a sample-compressed predictor. Theorem 2.1 can be used to obtain generalization bounds on the last task, which we do in mP2L when we compute the bound $\Psi$, but it holds for only one distribution of data. Therefore, Theorem 2.1 cannot be applied to bound all tasks learned by our continual learning scheme.
Le us now consider a reconstruction function that takes as input two compression sets $\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)$ and a message $\mu=(\mu_{1},\mu_{2})\in\mathscr{M}(\operatorname{\mathbf{j}},T)$. For any distribution $\operatorname{\mathcal{D}}_{t}$ and any dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on both compression sets as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}}(\theta)$ and ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{j}}}(\theta)$. The loss on the joint compression sets is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}\cup\operatorname{\mathbf{j}}}(\theta)$ and the loss on the complement set of the joint set $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}$ is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$ This formulation is important for the following theorem, as the reconstruction function should be fixed for all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
Theorem 3.1.
For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
$\forall t\in\{1,\ldots,T\},{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
$\leq\frac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\bigg{[}\log\matrixquantity(n_{t}\\ |{\mathbf{i}}|)+\log\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\mathbf{j}|)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i=1}^{T}\log\frac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\frac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\bigg{]}$
with $\mu=(\mu_{1},\mu_{2})$.
The proof is given in Appendix B. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting.
For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1. A more accurate predictor will achieve a tighter bound. However, a simpler predictor will achieve a tighter bound.
In the sample compression theory, we define the complexity of the model using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ is independent of its content but decays when $|\operatorname{\mathbf{j}}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets are, the smaller the probability of both messages will be and the complexity of the model grows when the compression sets grow.
In conclusion, the model that will achieve the best possible bound is a model that minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the smallest possible.
4 Experiments
Figure 1: Accuracy (above) and forgetting (below) over replay buffer size for (left column) MNIST dataset, (middle column) Fashion-MNIST dataset, (right column) EMNIST dataset. The orange triangles show the average performance for the corresponding box plot.
Figure 2: CoP2L bounds with respect to tasks on the MNIST and Fashion-MNIST datasets
Datasets We carry-out experiments on MNIST (LeCun et al., 2010), Fashion-MNIST (Xiao et al., 2017), and EMNIST (Cohen et al., 2017) datasets on class-incremental learning. We divide the tasks into increments of 2 classes. This yields 5 tasks for MNIST and Fashion-MNIST datasets, and 13 tasks for the EMNIST dataset.
Implementation Details We have used the Avalanche continual learning framework for our experimental setup (Carta et al., 2023). Avalanche is a popular toolkit for continual learning. We have used their class-incremental MNIST and Fashion-MNIST dataset implementations, as well as the replay baseline in order to ensure compatibility with the results in the literature. We implement a block version of mP2L, which adds $k$ examples to the compression set at a time, as described in Algorithm 2 of Paccagnan et al. (2024). For the experiments, we use $k=8$. We provide the code for our experimental results in an anonymous repository [^2] https://anonymous.4open.science/r/sample_compression_continual_learning-00C7..
Model architecture and training details We use a simple multilayer perceptron (MLP) architecture for all experiments (SimpleMLP class from the Avalanche toolkit), and train using the SGD optimizer, with a learning rate of 0.001 for CoP2L and 0.01 for Replay. For all experiments with the Replay method, we train the model 20 epochs on each tasks. In contrast, CoP2L follows a different training strategy, continuing until it reaches zero errors on the complement set. We add a weight of $\omega=15$ in the cross-entropy loss when training on previous tasks.
Computation of the bound To compute the bound, we need to keep track of the size of the compression sets and of the messages. We thus implement the compression function as described in Algorithm 4. However, we need to adjust the code to account for mP2L adding $k$ datapoints to the compression set at each iterations. When computing the message $\mu_{2}$, at iteration $t$, $\mu_{2}^{t}$ is the number of iterations of the mP2L algorithm. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by Viallard et al. (2021) to invert the $\mathrm{kl}$ divergence.
4.1 Metric Comparison over Buffer Size
In Figure 1, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks. Note that we denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as
$\text{Average Accuracy}_{T}=\frac{1}{s\cdot T}\sum_{i=1}^{s}\sum_{t=1}^{T}\text{Accuracy}_{T,t}^{i},$ (10)
where $i$ index is for different seeds, and $t$ is for the task index. In our experiments, we used $s=5$ seeds to measure the variability with respect to weight initialization. We also report the average forgetting at task $T$ which is defined as follows,
$\text{Average Forgetting}_{T}=$ (11)
$\frac{1}{(T-1)\cdot s}\sum_{t=1}^{T-1}\sum_{i=1}^{s}$ $\text{Accuracy}_{t,t}^{i}-\text{Accuracy}_{T,t}^{i},$
and can be calculated when $T>1$. Note that in Figure 1, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
We observe that for the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, we observe that for small buffer sizes, the average accuracy is obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better.
We would like to also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 1.
4.2 Empirical Study of the CoP2L generalization bound
In Figure 2, we showcase the generalization bound of CoP2L. We observe that the bound values for the generalization error always remain larger than the actual error incurred after each task, and we observe that the reported bound values are non-trivial. We also observe that the bound values generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 2, we provide the bounds for the models trained with a buffersize of 1000, but we give the results for all buffersizes, and all three datasets in Appendix C.
Figure 3: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function. Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
Figure 4: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right). We use 5 different task order, and show the variability with respect to each task.
4.3 Ablation on CoP2L loss terms
We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L algorithm (2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Figure 3, we show this experiment. We observe that for both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
4.4 Sensitivity analysis with respect to the task ordering
In order to also assess the sensitivity of CoP2L with respect to ordering of the tasks, on MNIST and Fashion-MNIST experiments we have conducted an experiment where we randomly shuffle the task order. For both dataset, we have tested our algorithm with 5 different random task ordering.
For each task, we report the cumulative accuracy
[…]
where $\text{Accuracy}_{T,t}^{o}$ denotes the accuracy obtained at task $t$ after finishing training until task $T$, and the $o$ index denotes the task order. We have used $O=5$ different task orderings. We present the results of this experiment in Figure 4. We observe that on both MNIST and Fashion-MNIST datasets CoP2L is able outperform with more tasks. We also observe that the variability of the results decrase more significantly for CoP2L compared to replay.
5 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they were non-vacuous. We furthermore showed that on three different datasets, our approach is able to outperform experience replay, which is known to be a strong continual learning baseline.
Acknowledgements
We also wish to thank Benjamin Leblanc for his help proof-reading the manuscript. Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192). Pascal Germain is supported by the NSERC Discovery grant RGPIN-2020-07223. Cem Subakan is supported by NSERC Discovery grant RGPIN-2023-05759. This research was also enabled in part by support provided by the Digital Research Alliance of Canada.
Appendix A Tighter sample compression $\mathrm{kl}$ bound
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound (Langford, 2005) for losses in $[0,1]$ (Foong et al., 2022).
Theorem A.1 ((Foong et al., 2022)).
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
$p\leq\mathrm{kl}^{-1}\quantity(\frac{1}{n}\sum_{i=1}^{n}X_{i},\frac{1}{n}\log\frac{1}{\delta}).$
We now prove a tighter $\mathrm{kl}$ sample compression bound. As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$. We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$. Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
Theorem A.2.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
Proof.
We prove the complement of the expression in Theorem A.2. We choose $\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$.
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (13)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (14)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (15)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$ (16)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$
$\leq\delta.$
We use the union bound in Equation 13 and Equation 14. In Equation 15, we use the i.i.d. assumption. In Equation 16 we use the Chernoff Test-set bound of Theorem A.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix B Proof of Theorem 3.1
We now prove Theorem 3.1.
See 3.1
Proof.
Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$. We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
Then, we have :
$\operatorname*{\mathbb{P}}_{S^{t}}\Bigg{(}\begin{split}&\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\\ &\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\end{split}\Biggr{)}$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem A.2. We can then lower bound this probability by $1-\delta$.
We finish the proof by applying the bound to all datasets.
With
$\epsilon(n_{k},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta},$
we have
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$\geq\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
$\geq 1-\delta$
We use a union bound argument to add the $\forall k\in[1,T]$, leading to the term $\frac{\delta}{T}$.
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of Marchand et al. (2003), which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$ and $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)$.
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $\quantity(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\nu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$. ∎
Appendix C Bound plots with different buffer sizes
In this section, we introduce the empirical bound estimates over tasks on MNIST, Fashion-MNIST and EMNIST datasets. In figure 5, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000).
Figure 5: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets