### v3

# Sample Compression for Self-Certified Continual Learning

Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees. In this paper, we present a new method called Continual Pick-to-Learn (CoP2L), which is able to retain the most representative samples for each task in an efficient way. CoP2L combines the Pick-to-Learn algorithm (rooted in the sample compression theory) and the experience replay continual learning scheme. This allows us to provide non-vacuous upper bounds on the generalization loss of the learned predictors, numerically computable after each task. We empirically evaluate our approach on several standard continual learning benchmarks across Class-Incremental, Task-Incremental, and Domain-Incremental settings. Our results show that CoP2L is highly competitive across all setups, often outperforming existing baselines, and significantly mitigating catastrophic forgetting compared to vanilla experience replay in the Class-Incremental setting. It is possible to leverage the bounds provided by CoP2L in practical scenarios to certify the predictor reliability on previously learned tasks, in order to improve the trustworthiness of the continual learning algorithm.
1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning de2021continual ; wang2023comprehensive , the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time. However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting McCloskey1989CatastrophicII ; forgetting_french ; goodfellow2013empirical . In order to cope with forgetting, various methodologies have been developed, such as regularization based approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architectural approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang , or rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der . It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting DellaLibera2023CLMASRAC ; Mai2020BatchlevelER ; chaudhry2019efficient ; aljundi2019onlinecontinuallearningmaximally . However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. paccagnan_pick_learn_2023 . P2L is a meta-algorithm that selects a compression set, which amounts to expressing the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory in mind floyd1995sample ; campi2023compression . By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor. This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning in the literature freund98selfbounding ; perezortiz2021tighter ; marks2025pick2learn-gp . This certification is arguably an asset to increase the trustworthiness towards machine learning systems.
In this paper, we propose Continual Pick-to-Learn (CoP2L), an algorithm that leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contributions can be summarized as follows:
(1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
(2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks, improving the trustworthiness of the continually learnt model.
(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain better or comparable performance compared with several strong continual learning baselines in most cases.
1.1 Related Work
Continual learning approaches.  In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories, which include regularization approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architecture-based approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang and rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
Continual learning generalization bounds.  In recent years, generalization bounds were proposed to study the asymptotic behavior of various continual learning schemes, for example using linear models Evron2022HowCC ; Goldfarb2022AnalysisOC ; pmlr-v202-lin23f , Neural Tangent Kernel (NTK) models trained with Orthogonal Gradient Descent AbbanaBennani2020GeneralisationGF ; Doan2020ATA ; farajtabar2019orthogonalgradientdescentcontinual , regularization-based methods yin2021optimizationgeneralizationregularizationbasedcontinual , or PackNet packnet_mallya inspired continual representation learning Li2022ProvableAE . Learning algorithms with convergence guarantees are were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features cao2022lifelong .
Different from all the listed works above, in this paper, we provide a generalization bound that is computed from the training set. Risk certificates on the test error are provided simultaneously for each task. Furthermore, also unlike the approaches above, our non-asymptotic bound is applicable to any neural network architecture, under minimal data assumptions, and does not have limitations on the continual learning setup that is employed.
Sample compression.  The sample compression theory has been shown effective for providing tight generalization bounds shah_margin-sparsity_2005 ; marchand2003set ; marchand2002set ; marchand2005learning ; shah_sample_2007 . However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks campi2023compression ; bazinet2024 and Gaussian Processes marks2025pick2learn-gp . In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
2 Background
Continual learning.  Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$. We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$. Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}(\theta_{T})\ \coloneqq\operatorname*{\mathbf{E}}_{({\mathbf{x}},y)\sim{\mathcal{D}}_{t}}\ell(\theta_{T},{\mathbf{x}},y)\,.$ (1)
The empirical loss counterparts on observed training samples are given by
^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting forgetting_french . Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
${\mathcal{F}}_{t}^{\ \text{replay}}(\theta_{t})\ \coloneqq\ \widehat{\mathcal{L}}_{S_{t}}(\theta_{t})+\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}\ell(\theta_{T},{\mathbf{x}}_{j},y_{j})\,,$ (3)
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as a baseline as it’s shown to be a simple and very effective continual learning strategy DellaLibera2023CLMASRAC ; Mai2020BatchlevelER .
Sample compression theory.  Given a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message. If this is the case, we call $f_{\theta}$ a sample-compressed predictor. The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that
$\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor. Although generally defined as a binary sequence, the message set can be defined by any set of countable sequences of symbols. Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 00 to $\infty$, constructed using the alphabet $\Sigma$. For all $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message. Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following common practices marchand2005learning ; bazinet2024 , we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})={n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k+1)^{-2}$. Also, given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
We now present a recent sample compression generalization guarantee bazinet2024 on which we build our theoretical framework. We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1 (bazinet2024 ).
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}{\in}\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}{\to}[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\mathbf{i}$ $\in\operatorname{\mathscr{P}}(n),\,\mu\in\mathscr{M}(\operatorname{\mathbf{i}}):$
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})])\,,$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
$\mathrm{kl}^{-1}(q,\varepsilon)=\mbox{\rm argsup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (5)
On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase. On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
Pick-to-Learn.  To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM boser_svm and the perceptron rosenblatt1958perceptron ; moran2020perceptron , have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) marchand2002set ; marchand2003set ; marchand2005learning and decision trees shah_sample_2007 , necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
Algorithm 1 Original Pick-To-Learn (P2L) paccagnan_pick_learn_2023
0:  $S=\{({\mathbf{x}}_{i},y_{i})\}_{i=1}^{n}$Training set}
0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
0:  $\gamma$Stopping criteria}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S}\ell(\theta_{0},{\mathbf{x}},y)$
3:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\geq\gamma$ do
4:    $k\leftarrow k+1$
5:    $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y})\}$
6:    $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
7:    $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)$
8:  end while
9:  return $\theta_{k}$Final parameters}
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan and Campi paccagnan_pick_learn_2023 as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm acts as an aditional training loop, by iteratively selecting datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of the original P2L in Algorithm 1. Recent works introduce improvements to the original scheme, notably early stopping strategies marks2025pick2learn-gp ; bazinet2024 and extension to the Bayesian setting marks2025pick2learn-gp .
3 Sample Compression for Continual Learning
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer to the compression set, thus mitigating the forgetting by focusing on a well-chosen datapoint.
Algorithm 2 Modified Pick-To-Learn (mP2L)
0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
0:  $S=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{n}$Training set (with weights)}
0:  $B^{\star}=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{m}$Buffer set (with weights)}
0:  $\gamma$Stopping criteria}
0:  $K^{\star}$Maximum number of iterations}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $S^{\star}\leftarrow S\cup B^{\star}$
3:  $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}}\ell(\theta_{0},{\mathbf{x}},y)\cdot w$
4:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\cdot\overline{w}\geq\gamma$ and $k\leq K^{\star}$ do
5:    $k\leftarrow k+1$
6:    $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y},\overline{w})\}$
7:    $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
8:    $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)\cdot w$
9:  end while
10:  if $k<K^{\star}$ then
11:    $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$Eq. (6)}
12:  end if
13:  return $\theta_{k},C_{k}$. {Final parameters and compression set}
Algorithm 3 Continual Pick-To-Learn (CoP2L)
0:  $\theta_{0}${Initial parameters of the model}
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
0:  $\gamma$mP2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
1:  $B_{i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:    $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\ \forall i{=}1,\ldots,t{-}1$
7:     $B_{t}\leftarrow\mbox{sample}(\hat{S}_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
8:    $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
9:  end for
10:  return $\theta_{T}$. {Final parameters}
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm (Algorithm 3, entitled CoP2L) calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
We modified Pick-To-Learn in two significant ways. (i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer. (ii) The second modification is the choice of the returned predictor by mP2L. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior, P2L adding most training samples into its compression set, which in turn gives a large generalization bound value. Instead, we perform early stopping based on the bound value (as proposed by bazinet2024 ). That is, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
$\Psi(S,\theta,C)$ $=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}\quantity(\theta),\ \tfrac{1}{|S\setminus C|}\log\quantity(\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta})),$ (6)
with $\mathrm{kl}^{-1}$ defined by Equation 5. Thanks to Theorem 2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with probability at least $1-\delta$.
Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L). Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$. It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks. The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
Compression and reconstruction scheme.  From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample$(S,m)$ in CoP2L cannot be reproduced in the reconstruction step. In Appendix A, we provide the compression and reconstruction functions of CoP2L, which give rise to sample compression bounds as presented in Theorem 3.1. Noteworthy, after learning on $T$ tasks, the compression function provides two sample compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as $\mathscr{M}(\operatorname{\mathbf{j}},T)=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ As explained further in appendix, the use of two compression sets, along with the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves to the proper treatement of the buffer $B^{\star}$, while the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
Generalization bound.  Given the compression and reconstruction function detailed in Apprendix A, Theorem 2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$. However, this bound holds for only one distribution of data. Therefore, it cannot be applied to bound the risks of the learned predictor on previously tasks seen by CoP2L. The next Theorem 3.1 holds for any previously learned tasks. For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $(S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}})^{\rm\bf c}$ as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\big{(}S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big{)}.$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
Theorem 3.1.
For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
$\forall t\in\{1,\ldots,T\},{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu=(\mu_{1},[\mu_{2}^{i}]_{i=1}^{T})\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
$\leq\tfrac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\big{[}\log{\textstyle{n_{t}\choose|{\mathbf{i}}|}}+\log{\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}}+\textstyle\sum_{i=1}^{T}\log\tfrac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\tfrac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\big{]}\,.$
The proof is given in Appendix C. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting. For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1 between the predictor accuracy and its complexity. Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ decays when $|\mathbf{j}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets, the smaller the probability of both messages. In conclusion, the bound encodes the ability of a model to minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the small.
4 Experiments
Datasets and continual learning settings.  We carry-out experiments on MNIST lecun2010mnist , Fashion-MNIST fashionmnist , EMNIST (characters) cohen2017emnistextensionmnisthandwritten , PermutedMNIST goodfellow2013empirical , Rotated MNIST, CIFAR10 cifar10cifar100 , CIFAR100 cifar10cifar100 and TinyImageNet tiny-imagenet . We report results that include Class-Incremental (CI), Task-Incremental (TI), and Domain-Incremental (DI) continual learning settings. The dataset stats regarding how many classes/data items are available per task is given in Appendix H.
Baselines.  To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental, Task-Incremental, and Domain-Incremental continual learning settings. For each continual learning set-up we include baselines that yield the best performance in each category according to the Avalanche benchmark lomonaco2021avalanche .
For the Class-Incremental setting, we use the baselines Replay rolnick2019experience , iCaRL rebuffi2017icarl , and GDumb prabhu2020gdumb . Replay in particular is considered a strong baseline DellaLibera2023CLMASRAC due to its simplicity and effectiveness. In the Task-Incremental setting, we compare against Replay rolnick2019experience , LaMAML gupta2020look and Learning without Forgetting (LwF) li2017learning . These methods are well suited for TI scenarios where task identity is known at test time. In the Domain-Incremental setting, we evaluate against Replay rolnick2019experience , Less Forgetful Learning (LFL) jung2018less , Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , and Synaptic Intelligence (SI) zenke2017continual .
Implementation.  All our experiments are conducted using the Avalanche continual learning framework avalanchelib , which provides a unified interface for datasets, benchmarks, and baseline implementations. We also use the official ContinualAI benchmark repository [^1] https://github.com/ContinualAI/continual-learning-baselines for consistent implementation and fair comparisons. The code for the experimental results is available in the supplemental material.
Model architecture and training details.  We use the SimpleMLP architecture and customized SimpleCNN architectures (SimpleMLP and SimpleCNN from the Avalanche toolkit) for the MNIST, FMNIST, and EMNIST experiments. All models are trained using the SGD optimizer. The details of the models architectures and the hyperparameters used can be found in Appendix G. For experiments on more complex datasets such as CIFAR10, CIFAR100, and TinyImageNet, we use the pretrained feature extractor Dino caron2021emerging using ViT dosovitskiy2020image and ResNet50 he2016deep backbones, followed by a trainable MLP head. This choice is motivated by the insight that CoP2L can further benefit from stronger feature representations, especially in settings with limited data per class or high dataset complexity. By leveraging Dino’s obtained embeddings, we enhance the quality of the input representations, which we empirically observe to yield good performance. Results with standard deviations across seeds and total training times are reported in Appendix F. We used a single NVIDIA 4090 GPU for our experiments.
Accuracy and Forgetting.  We denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as,
. $\textit{Average Accuracy}_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}_{T,t}$, where $t$ is for the task index. .
We also report the average forgetting at task $T$ which is defined as
$\textit{Average Forgetting}_{T}=\frac{1}{(T-1)}\sum_{t=1}^{T-1}[\text{Accuracy}_{t,t}-\text{Accuracy}_{T,t}],$ which is calculated when $T>1$.
4.1 Main Results
In Table 1, we present the average accuracy and average forgetting values defined in the section above, obtained on MNIST, FMNIST and EMNIST dataset, both with an MLP and a CNN architecture, where the networks are trained from scratch. In Appendix F, we also provide the standard deviations as well. We observe that with the MLP architecture CoP2L is able to outperform the baselines. Furthermore, we also observe that CoP2L is able to obtain competitive results when employed with a CNN architecture as well. In Appendix E, we also provide an analysis of the performance of CoP2L for varying buffer sizes.
In Table 2, we extended our results to cover CIFAR10, CIFAR100, and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings. We also provide results for Domain-Incremental setting with Permuted-MNIST goodfellow2013empirical and Rotated-MNIST datasets.
Table 1: Summary of average accuracies and forgetting values obtained on MNIST, FMNIST, EMNIST (5 tasks, Class-Incremental), and on Permuted-MNIST and Rotated-MNIST (10 tasks, Domain-Incremental).
Method / Dataset MNIST FMNIST EMNIST
Acc / For Acc / For Acc / For
CoP2L (MLP) 95.1 / 3.4 84.2 / 12.1 78.0 / 16.7
Replay (MLP) 94.5 / 6.1 84.0 / 15.9 77.2 / 22.1
GDumb (MLP) 91.7 / 3.8 81.8 / 10.4 62.6 / 16.0
CoP2L (CNN) 96.6 / 2.2 87.4 / 11.1 87.9 / 8.4
Replay (CNN) 96.9 / 3.3 87.6 / 13.3 87.7 / 12.3
GDumb (CNN) 94.4 / 3.0 85.0 / 8.8 85.1 / 7.5
Method / Dataset PermMNIST RotMNIST
Acc / For Acc / For
CoP2L (MLP) 92.3 / 6.0 88.3 / 9.8
Replay (MLP) 91.5 / 6.4 88.2 / 10.0
LFL (MLP) 83.5 / 9.3 51.6 / 45.7
EWC (MLP) 83.7 / 5.2 56.5 / 35.4
SI (MLP) 86.7 / 10.7 57.7 / 43.9
Table 2: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets. CI denotes Class-Incremental, TI denotes Task-Incremental continual learning settings.
CIFAR10 (5 tasks), CI CIFAR100 (10 tasks), CI CIFAR100 (20 tasks), CI TinyImNet (40 tasks), CI
Method Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$)
CoP2L-Dino (ViT) 94.45 3.36 73.83 16.76 72.50 16.96 51.13 36.81
Replay-Dino (ViT) 94.01 6.01 68.51 30.44 69.23 29.60 47.66 48.15
iCaRL-Dino (ViT) 93.24 1.55 61.50 14.53 64.11 15.25 20.78 52.95
GDumb-Dino (ViT) 94.15 3.39 72.73 11.70 72.05 11.38 54.53 14.79
CoP2L-Dino (RN50) 81.13 5.55 49.18 21.99 49.98 24.32 32.74 36.40
Replay-Dino (RN50) 81.83 17.09 49.40 42.07 49.91 44.41 28.43 63.26
iCaRL-Dino (RN50) 74.56 1.54 45.97 8.71 46.87 6.32 33.13 10.95
GDumb-Dino (RN50) 80.70 8.95 43.83 20.31 43.78 19.03 24.04 21.82
CIFAR10 (5 tasks), TI CIFAR100 (10 tasks), TI CIFAR100 (20 tasks), TI TinyImNet (40 tasks), TI
CoP2L-Dino (ViT) 99.01 -0.03 95.15 0.92 97.10 0.71 94.28 1.22
Replay-Dino (ViT) 99.24 -0.03 96.59 0.23 98.06 0.41 95.82 0.36
LaMAML-Dino (ViT) 99.18 0.06 95.77 0.40 97.70 0.33 94.90 1.15
LwF-Dino (ViT) 97.53 2.08 95.26 1.06 95.73 2.15 85.79 9.09
CoP2L-Dino (RN50) 96.17 0.79 86.58 2.26 90.88 2.37 88.75 2.13
Replay-Dino (RN50) 96.96 0.42 87.59 0.64 93.13 0.16 90.53 0.74
LaMAML-Dino (RN50) 96.71 0.66 83.13 6.81 90.57 3.35 87.25 4.55
LwF-Dino (RN50) 95.90 0.15 79.29 0.67 87.94 0.09 85.15 0.90
Figure 1: Example demonstration of emprical upper bounds on test errors, on Class-Incremental (top-left CIFAR10, top-right CIFAR100) and Task-Incremental (bottom-left CIFAR10, bottom-right CIFAR100) datasets.
Empirical study of the CoP2L generalization bound.  In Figure 1, we showcase the empirical upper bounds on test errors. The details on the computation of the bound can be found in the Appendix  A. We observe that the bound values are non-trivial and generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 1, we provide plots for the bounds for the models trained with a buffer size of 2000, for Class-Incremental and Task-Incremental settings. CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task. These conditions allow the model to better compress task-specific knowledge, leading to both stronger theoretical guarantees and improved empirical performance. We also observe that with the ViT backbone, the bounds are significantly tighter than those obtained with the ResNet50 backbone. This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L.
We also give the results for all buffer sizes for the three MNIST datasets, and also bound estimates for the CIFAR10, CIFAR100, and ImageNet datasets (TinyImageNet tiny-imagenet , ImageNette imagenette and ImageWoof Howard_Imagewoof_2019 ) as well as bounds for the Domain-Incremental setting in Appendix D.
4.2 Analysis of CoP2L
In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size. Due to space limitations the figures are provided in the appendix, however the main results are mentioned here in the text.
Ablation on CoP2L loss terms  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Appendix E, in Figure 10, we show this experiment. For both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
[…]
Analysis on buffer size.  In Appendix E, in Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
For the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, when using small buffer sizes, the average accuracy obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better. We also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
5 Limitations and Broader Impact Statement
We observe with our experiments that the performance of CoP2L is dependent to the ability of the feature extractor (backbone). In case a weak backbone is employed, CoP2L could necessitate a larger compression set (e.g. more data points per task). We have also seen that in those cases CoP2L might take longer to converge. We noted in our experiments that using deep pretrained networks such as ViT results in good performance, by increasing the compression ability. More broadly, we do not believe there is any particular negative social impact to note for our proposed methodology. On the contrary, we believe that our algorithm could be instrumental for improving trustworthiness of continual learning / lifelong learning in practice.
6 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory for self-certified continual learning. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative. We furthermore showed that on several challenging datasets, that our approach yields comparable or better results compared several strong baselines, while also providing learning guarantees, for trustworthy continual learning. As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could also lead to new compelling ways of obtaining self-certified predictors.
Acknowledgements
Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192). Pascal Germain, and Cem Subakan are supported by Discovery Grants RGPIN-2020-07223 and RGPIN 2023-05759, respectively. This research was also enabled in part by support provided by Calcul Québec (https://www.calculquebec.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/en).
Appendix A Compression and Reconstruction Functions
In line with the prevalent literature, we introduced the sample compression framework (Section 2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^2] The idea of using multiple compression sets appeared in [40] and [39], but in a setting without a message.
Algorithm 4 Compression function of Continual Pick-To-Learn (CoP2L)
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
1:  $B_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{i}}}_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{j}}}_{t}\leftarrow\emptyset,\mu_{1,t}\leftarrow\emptyset\quad\forall t=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset,\mu_{2}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:    $\mu_{2}^{t}\leftarrow|C^{\star}|$
7:    $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
8:    $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
9:    $B_{t}\leftarrow\mbox{sample}(S_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
10:    for $i\in\{1,\ldots,t-1\}$ do
11:       for $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\in B^{\star}$ and $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\notin B_{i}$ do
12:          $S^{\operatorname{\mathbf{j}}}_{i}\leftarrow S^{\operatorname{\mathbf{j}}}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
13:          $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
14:          $\mu_{1,i}^{j}\leftarrow\{t\}$
15:       end for
16:    end for
17:    $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
18:  end for
19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$. {Learned parameters, compression sets and messages sets}
A.1 Compression Function
Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function. In this subsection, we expose each of these challenges while explaining how we the proposed compression algorithm (Algorithm 4) allows to circumvent them.
The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3). When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second challenge is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
After learning on $T$ tasks, the compression function provides the sample compression set $S^{\operatorname{\mathbf{i}}}$ and  $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (7)
A.2 Reconstruction Function
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is addressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
Algorithm 5 Reconstruction function of Continual Pick-To-Learn (CoP2L)
0:  $S^{\operatorname{\mathbf{i}}}_{1},S^{\operatorname{\mathbf{i}}}_{2},\ldots,S^{\operatorname{\mathbf{i}}}_{T}$Compression sets}
0:  $S^{\operatorname{\mathbf{j}}}_{1},S^{\operatorname{\mathbf{j}}}_{2},\ldots,S^{\operatorname{\mathbf{j}}}_{T}$Compression sets}
0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T}${Message sets}
0:  $\mu_{2}${Message set}
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
1:  $B^{\star}\leftarrow\emptyset$
2:  for $t\in\{1,\ldots,T\}$ do
3:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S^{\operatorname{\mathbf{i}}}_{t}}$
4:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\mu_{2}^{t})$
5:    $B_{t}\leftarrow(S^{\operatorname{\mathbf{i}}}_{t}\setminus C^{\star})\cup S^{\operatorname{\mathbf{j}}}_{t}$
6:    for $i\in\{1,\ldots,t-1\}$ do
7:       for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in S^{\operatorname{\mathbf{j}}}_{i}$ do
8:          if $\mu_{1,i}^{j}=t$ then
9:             $B_{i}\leftarrow B_{i}\setminus\{(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\}$
10:          end if
11:       end for
12:       $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
13:    end for
14:  end for
15:  return $\theta_{T}$. {Learned parameters}
A.3 Computation of the Bound
To compute the bound of Theorem 3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$. By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually. Moreover, we are completely unable to retrieve the messages. Thus, instead of implementing Algorithm 3, we actually implemented the compression function as described in Algorithm 4.
Moreover, we use a block version of P2L, as defined by Algorithm 2 of [23]. We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Appendix G. Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm 4) becomes the number of iterations of the mP2L algorithm at for task $t$. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by [65] to invert the $\mathrm{kl}$ divergence.
Appendix B Tighter sample compression $\mathrm{kl}$ bound
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound [66] for losses in $[0,1]$ [67].
Theorem B.1 ([67]).
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
$p\leq\mathrm{kl}^{-1}\quantity(\tfrac{1}{n}\textstyle\sum_{i=1}^{n}X_{i},\tfrac{1}{n}\log\tfrac{1}{\delta}).$
We now prove a tighter $\mathrm{kl}$ sample compression bound. As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$. We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$. Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
Theorem B.2.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
Proof.
We prove the complement of the expression in Theorem B.2. Denote the upper bound
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (8)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (9)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (10)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$ (11)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$
$\leq\delta.$
We use the union bound in Equation 8 and Equation 9. In Equation 10, we use the i.i.d. assumption. In Equation 11 we use the Chernoff Test-set bound of Theorem B.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix C Proof of Theorem 3.1
See 3.1
Proof.
Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$. We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
Let
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
Then, we have :
$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem B.2. We can then lower bound this probability by $1-\delta$.
We finish the proof by applying the bound to all datasets. We have
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$\geq$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
$\geq$ $1-\delta.$
We use a union bound argument to add the $\forall t\in[1,T]$, leading to the term $\frac{\delta}{T}$.
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of [39], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
$P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\quad\mbox{and }\quad P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)\,.$
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$ (Equation 7). ∎
Appendix D Bound Plots
D.1 Bound plots for different buffer sizes
In this section, we present the computed bound values over task risks on MNIST, Fashion-MNIST and EMNIST datasets (letters). In Figure 2, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000), obtained using the MLP architecture. We also show the results obtained with the CNN architecture in Figure 3. We report the estimated deviation of the bound values over different seeds using the shaded regions which show the $\pm$1 standard deviation, and also the whiskers (plotted using the seaborn toolkit’s boxplot function) which show the spread of the values excluding the outliers. This plotting style for showing the deviation across seeds applies to all bound plots presented in this paper.
Figure 2: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using an MLP architecture
Figure 3: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using a CNN architecture
We observe from Figures 2 and 3 that the MLP architecture and the CNN architecture give non-trivial bounds that follow the trends exhibited by the test error curve.
D.2 Bound plots for CIFAR10, CIFAR100, TinyImageNet and ImageNet (subsets) datasets
In this section, we present the plots for the bound values obtained for the additional datasets, namely CIFAR10, CIFAR100, and Imagenet subsets. The Imagenet subsets include the Tinyimagenet [51], ImageWoof [64], and ImageNette [63] datasets. We provide the results on ImageWoof and ImageNette to showcase that the estimated bounds get tighter when more data is avaible per each task.
Figure 4: CoP2L bounds with respect to tasks on the CIFAR10 (5 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
Figure 5: CoP2L bounds with respect to tasks on the CIFAR100 (10 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
Figure 6: CoP2L bounds with respect to tasks on the CIFAR100 (20 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
Figure 7: CoP2L bounds with respect to tasks on the TinyImageNet (40 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
Figure 8: CoP2L bounds with respect to tasks on the ImageNette (5 tasks) dataset (left) and ImageWoof dataset (right) in Task-Incremental setting
Figure 9: CoP2L bounds with respect to tasks on the PermutedMNIST (10 tasks) dataset (left) and on the RotatedMNIST dataset in Domain-Incremental setting
Appendix E Analysis of CoP2L
In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size. The text matches the one provided in the main paper, but we also include the aforementioned figures.
Ablation on CoP2L loss terms.  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Figure 10, we show the results of this experiment. For both accuracy (left panel) and forgetting (right panel), we observe that the inclusion of the additional loss weighting term significantly improves the results.
[…]
Figure 10: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function. Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
Figure 11: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right). We use 5 different task order, and show the variability with respect to each task.
E.1 Analysis with respect to buffersize
In Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
Note that in Figure 12, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
We observe that for the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, we observe that for small buffer sizes, the average accuracy is obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better.
We would like to also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
Figure 12: Accuracy (left) and forgetting (right) over replay buffer size for (top row) MNIST dataset, (middle row) Fashion-MNIST dataset, (bottom row) EMNIST dataset. The orange triangles show the average performance for the corresponding box plot.
Appendix F All results with standard deviation and training time measurement
In this section we provide the average accuracy values obtained on MNIST, FMNIST, EMNIST, CIFAR10, CIFAR100, TinyImagenet, PermutedMNIST and RotatedMNIST datasets with the standard deviation estimates. Note that Table 2 of the main paper provides the results obtained with a single seed, whereas the following tables provide the average results over five seeds, and the corresponding standard deviations. Note that we use different random number generator seeds to account for variability in random number generations (e.g. model parameter initializations) – but train / test splits are kept fixed with respect to seeds across different experiments.
Note that, for CIFAR10, CIFAR100, TinyImagenet, ImageWoof, ImageNette, PermutedMNIST and RotatedMNIST experiments we also report the total time it takes to complete each experiment (for all tasks).
CIFAR10 (5 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $94.45\pm 0.14$ $2.10\pm 0.28$ $118.07\pm 2.59$
Replay-Dino (ViT) $94.00\pm 0.19$ $6.11\pm 0.25$ $486.25\pm 43.07$
iCaRL-Dino (ViT) $88.60\pm 10.04$ $3.13\pm 4.20$ $884.59\pm 1.39$
GDumb-Dino (ViT) $94.16\pm 0.24$ $3.16\pm 0.39$ $47.72\pm 0.12$
CoP2L-Dino (ResNet50) $80.98\pm 0.17$ $5.84\pm 0.37$ $527.95\pm 8.43$
Replay-Dino (ResNet50) $81.38\pm 0.39$ $17.72\pm 0.51$ $287.38\pm 25.06$
iCaRL-Dino (ResNet50) $74.49\pm 0.22$ $1.79\pm 0.43$ $579.17\pm 2.39$
GDumb-Dino (ResNet50) $80.84\pm 0.14$ $9.35\pm 0.30$ $31.98\pm 0.12$
CIFAR100 (10 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $73.49\pm 0.27$ $17.06\pm 0.29$ $522.18\pm 3.90$
Replay-Dino (ViT) $68.71\pm 0.25$ $30.02\pm 0.31$ $453.04\pm 11.93$
iCaRL-Dino (ViT) $65.54\pm 2.23$ $13.64\pm 0.64$ $2491.81\pm 2889.82$
GDumb-Dino (ViT) $72.51\pm 0.26$ $11.53\pm 0.65$ $95.43\pm 0.34$
CoP2L-Dino (ResNet50) $49.53\pm 0.24$ $22.31\pm 0.20$ $1724.15\pm 23.50$
Replay-Dino (ResNet50) $49.41\pm 0.38$ $42.12\pm 0.54$ $295.80\pm 15.88$
iCaRL-Dino (ResNet50) $46.06\pm 0.40$ $8.50\pm 0.28$ $676.76\pm 56.25$
GDumb-Dino (ResNet50) $43.37\pm 0.24$ $20.23\pm 0.42$ $63.58\pm 0.36$
CIFAR100 (20 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $70.56\pm 0.99$ $21.15\pm 2.10$ $820.66\pm 176.27$
Replay-Dino (ViT) $69.21\pm 0.26$ $29.52\pm 0.25$ $457.46\pm 2.07$
iCaRL-Dino (ViT) $62.36\pm 1.46$ $15.88\pm 0.84$ $1359.13\pm 4.58$
GDumb-Dino (ViT) $72.35\pm 0.30$ $11.77\pm 0.40$ $190.80\pm 0.42$
CoP2L-Dino (ResNet50) $46.74\pm 1.63$ $22.54\pm 0.92$ $1771.12\pm 757.50$
Replay-Dino (ResNet50) $49.44\pm 0.30$ $44.96\pm 0.36$ $293.25\pm 1.77$
iCaRL-Dino (ResNet50) $46.47\pm 0.33$ $6.74\pm 0.31$ $858.11\pm 120.41$
GDumb-Dino (ResNet50) $43.60\pm 0.38$ $19.93\pm 0.56$ $126.74\pm 0.46$
TinyImageNet (40 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $51.59\pm 0.24$ $36.32\pm 0.29$ $1451.93\pm 4.71$
Replay-Dino (ViT) $47.49\pm 0.47$ $48.39\pm 0.49$ $1128.97\pm 16.95$
iCaRL-Dino (ViT) $21.06\pm 1.11$ $34.49\pm 9.37$ $3919.23\pm 2170.08$
GDumb-Dino (ViT) $53.93\pm 0.34$ $15.11\pm 0.19$ $471.23\pm 17.85$
CoP2L-Dino (ResNet50) $33.30\pm 0.32$ $35.93\pm 0.27$ $3671.69\pm 29.19$
Replay-Dino (ResNet50) $28.38\pm 0.42$ $63.36\pm 0.44$ $882.12\pm 180.18$
iCaRL-Dino (ResNet50) $33.89\pm 0.47$ $10.42\pm 0.44$ $1905.26\pm 295.18$
GDumb-Dino (ResNet50) $24.29\pm 0.58$ $21.29\pm 0.56$ $332.22\pm 0.98$
CIFAR10 (5 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $99.04\pm 0.05$ $-0.00\pm 0.10$ $64.51\pm 0.68$
Replay-Dino (ViT) $99.29\pm 0.04$ $-0.01\pm 0.03$ $217.83\pm 0.76$
LaMAML-Dino (ViT) $99.25\pm 0.06$ $0.05\pm 0.07$ $344.65\pm 1.46$
LwF-Dino (ViT) $97.89\pm 0.41$ $1.69\pm 0.48$ $165.41\pm 0.75$
CoP2L-Dino (ResNet50) $96.22\pm 0.18$ $0.56\pm 0.28$ $172.51\pm 4.42$
Replay-Dino (ResNet50) $96.88\pm 0.06$ $0.42\pm 0.13$ $144.57\pm 0.51$
LaMAML-Dino (ResNet50) $96.65\pm 0.21$ $0.73\pm 0.21$ $310.51\pm 1.64$
LwF-Dino (ResNet50) $95.80\pm 0.15$ $0.33\pm 0.20$ $97.09\pm 0.21$
CIFAR100 (10 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $95.03\pm 0.08$ $0.90\pm 0.11$ $290.95\pm 30.98$
Replay-Dino (ViT) $96.43\pm 0.14$ $0.34\pm 0.17$ $235.81\pm 1.10$
LaMAML-Dino (ViT) $95.58\pm 0.16$ $0.58\pm 0.12$ $397.78\pm 3.40$
LwF-Dino (ViT) $95.09\pm 0.10$ $0.98\pm 0.10$ $173.57\pm 0.63$
CoP2L-Dino (ResNet50) $86.78\pm 0.16$ $2.01\pm 0.17$ $435.10\pm 7.74$
Replay-Dino (ResNet50) $87.61\pm 0.05$ $0.86\pm 0.16$ $157.05\pm 0.14$
LaMAML-Dino (ResNet50) $83.57\pm 0.43$ $6.18\pm 0.53$ $357.24\pm 4.58$
LwF-Dino (ResNet50) $79.39\pm 0.13$ $0.47\pm 0.11$ $101.27\pm 0.48$
CIFAR100 (20 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $97.18\pm 0.13$ $0.56\pm 0.18$ $226.65\pm 21.43$
Replay-Dino (ViT) $98.09\pm 0.09$ $0.25\pm 0.12$ $250.78\pm 1.45$
LaMAML-Dino (ViT) $97.84\pm 0.10$ $0.26\pm 0.14$ $448.43\pm 2.81$
LwF-Dino (ViT) $95.74\pm 0.19$ $2.15\pm 0.21$ $180.36\pm 0.98$
CoP2L-Dino (ResNet50) $90.89\pm 0.22$ $2.27\pm 0.28$ $278.67\pm 6.20$
Replay-Dino (ResNet50) $92.89\pm 0.18$ $0.47\pm 0.24$ $167.65\pm 1.26$
LaMAML-Dino (ResNet50) $89.89\pm 0.75$ $4.12\pm 0.78$ $407.84\pm 7.38$
LwF-Dino (ResNet50) $87.73\pm 0.18$ $0.33\pm 0.23$ $105.46\pm 0.49$
TinyImageNet (40 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $93.81\pm 0.26$ $1.50\pm 0.28$ $1039.14\pm 146.21$
Replay-Dino (ViT) $95.66\pm 0.08$ $0.50\pm 0.09$ $632.18\pm 2.97$
LaMAML-Dino (ViT) $94.82\pm 0.23$ $0.79\pm 0.19$ $1527.24\pm 891.43$
LwF-Dino (ViT) $85.18\pm 0.76$ $9.72\pm 0.71$ $430.63\pm 2.14$
CoP2L-Dino (ResNet50) $88.57\pm 0.17$ $2.46\pm 0.19$ $890.94\pm 3.44$
Replay-Dino (ResNet50) $90.03\pm 0.26$ $1.26\pm 0.27$ $514.42\pm 102.82$
LaMAML-Dino (ResNet50) $87.20\pm 0.34$ $4.53\pm 0.28$ $1268.50\pm 551.93$
LwF-Dino (ResNet50) $85.03\pm 0.17$ $0.89\pm 0.21$ $284.79\pm 5.67$
ImageNette (5 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $98.22\pm 0.19$ $-0.43\pm 0.12$ $118.71\pm 24.35$
ImageWoof (5 tasks) Task-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L-Dino (ViT) $93.34\pm 0.30$ $-0.65\pm 0.40$ $246.77\pm 13.43$
Permuted-MNIST (10 tasks) Domain-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L $91.61\pm 0.39$ $6.69\pm 0.40$ $1381.06\pm 1232.99$
Replay $91.28\pm 0.18$ $6.75\pm 0.21$ $1128.99\pm 49.82$
LFL $82.76\pm 1.23$ $10.15\pm 1.36$ $92.86\pm 0.24$
EWC $82.53\pm 0.82$ $6.50\pm 0.88$ $340.16\pm 3.81$
SI $85.85\pm 0.61$ $11.69\pm 0.67$ $319.95\pm 1.38$
Rotated-MNIST (10 tasks) Domain-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
CoP2L $89.40\pm 1.00$ $8.71\pm 1.04$ $4970.86\pm 1452.34$
Replay $88.75\pm 0.67$ $9.53\pm 0.72$ $3894.62\pm 75.11$
LFL $53.09\pm 7.06$ $44.18\pm 7.93$ $387.61\pm 160.54$
EWC $55.59\pm 5.57$ $36.58\pm 6.27$ $1133.19\pm 18.10$
SI $59.90\pm 6.48$ $41.38\pm 7.18$ $1043.86\pm 7.28$
MNIST (5 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%)
CoP2L (MLP) $95.12\pm 0.18$ $3.35\pm 0.24$
Replay (MLP) $94.48\pm 0.28$ $6.08\pm 0.32$
GDumb (MLP) $91.70\pm 0.31$ $3.84\pm 0.29$
CoP2L (CNN) $96.61\pm 0.07$ $2.16\pm 0.28$
Replay (CNN) $96.93\pm 0.29$ $3.31\pm 0.38$
GDumb (CNN) $94.41\pm 0.13$ $2.97\pm 0.46$
Fashion-MNIST (5 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%)
CoP2L (MLP) $84.23\pm 0.27$ $12.10\pm 0.66$
Replay (MLP) $83.99\pm 0.52$ $15.92\pm 0.42$
GDumb (MLP) $81.76\pm 0.36$ $10.44\pm 0.50$
CoP2L (CNN) $87.35\pm 0.15$ $11.12\pm 0.36$
Replay (CNN) $87.59\pm 0.11$ $13.32\pm 0.29$
GDumb (CNN) $84.98\pm 0.21$ $8.78\pm 0.41$
EMNIST (13 tasks) Class-Incremental
Methods Stream Accuracy (%) Stream Forgetting (%)
CoP2L (MLP) $78.03\pm 0.17$ $16.67\pm 0.31$
Replay (MLP) $77.22\pm 0.44$ $22.08\pm 0.50$
GDumb (MLP) $62.64\pm 0.20$ $15.96\pm 0.58$
CoP2L (CNN) $87.90\pm 0.41$ $8.38\pm 0.54$
Replay (CNN) $87.70\pm 0.50$ $12.29\pm 0.49$
GDumb (CNN) $85.05\pm 0.27$ $7.50\pm 0.61$
Appendix G Training, architecture details and hyperparameter settings for the experiments
[…]
In experiments using a Dino backbone (both ResNet50 and ViT backbones), we upsampled the images to 64x64, since Dino was pretrained on 224x224 images. We observed better performance from the pretrained feature extractor when the input image size is closer to the original training image size of Dino.
We present below the two custom CNN architectures derive from the SimpleCNN architecture (from the Avalanche toolkit). All other architectures used for the experiments are available in the Avalanche toolkit.
The specific CNN architecture used for MNIST experiments is defined as follows:
MnistCnn(
  (cnn_layers): Sequential(
    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
    (1): ReLU()
    (2): Dropout2d(p=0.2, inplace=False)
    (3): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
    (4): ReLU()
    (5): Dropout2d(p=0.2, inplace=False)
    (6): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
  (linear_layers): Sequential(
    (0): Linear(in_features=9216, out_features=128, bias=True)
    (1): ReLU()
    (2): Dropout(p=0.2, inplace=False)
    (3): Linear(in_features=128, out_features=10, bias=True)
The specific CNN architecture used for FMNIST and EMNIST experiments is defined as follows:
EMNISTFashionMnistCnn(
  (cnn_layers): Sequential(
    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
    (1): ReLU()
    (2): BatchNorm2d(32, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
    (3): Dropout2d(p=0.3, inplace=False)
    (4): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
    (5): ReLU()
    (6): BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
    (7): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
    (8): Dropout2d(p=0.3, inplace=False)
    (9): Conv2d(64, 128, kernel_size=(3, 3), stride=(1, 1))
    (10): ReLU()
    (11): BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
    (12): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
    (13): Dropout2d(p=0.3, inplace=False)
  (linear_layers): Sequential(
    (0): Linear(in_features=3200, out_features=256, bias=True)
    (1): ReLU()
    (2): Dropout(p=0.3, inplace=False)
    (3): Linear(in_features=256, out_features=10, bias=True)
Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
CoP2L 8 15.0 10 256 0.001
Replay NA NA 20 128 0.01
GDumb NA NA 20 128 0.01
Table 3: Hyperparameter setup for MNIST, FMNIST and EMNIST using a MLP architecture. NA signifies Not Applicable.
Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
CoP2L (MNIST & FMNIST) 48 25.0 25 256 0.0009
CoP2L (EMNIST) 105 20.0 15 130 0.0006
Replay NA NA 20 128 0.01
GDumb NA NA 20 128 0.01
Table 4: Hyperparameter setup for MNIST, FMNIST and EMNIST using a CNN architecture. NA signifies Not Applicable.
Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
CoP2L-Dino (ViT) 4 25.0 2 256 0.001
Replay-Dino (ViT) NA NA 20 128 0.01
iCaRL-Dino (ViT) NA NA 10 128 0.01
GDumb-Dino (ViT) NA NA 20 128 0.01
CoP2L-Dino (RN50) 16 25.0 2 256 0.001
Replay-Dino (RN50) NA NA 20 128 0.01
iCaRL-Dino (RN50) NA NA 30 128 0.01
GDumb-Dino (RN50) NA NA 20 128 0.01
Table 5: Hyperparameter setup for Class-Incremental setup. NA signifies Not Applicable.
Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
CoP2L-Dino (ViT) 4 1.0 2 256 0.001
Replay-Dino (ViT) NA NA 10 128 0.01
LaMAML-Dino (ViT) NA NA 10 10 0.01
LwF-Dino (ViT) NA NA 10 200 0.001
CoP2L-Dino (RN50) 16 1.0 2 256 0.001
Replay-Dino (RN50) NA NA 10 128 0.01
LaMAML-Dino (RN50) NA NA 10 10 0.01
LwF-Dino (RN50) NA NA 10 200 0.001
Table 6: Hyperparameter setup for Task-Incremental setup. NA signifies Not Applicable.
Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
CoP2L 16 15.0 1 256 0.001
Replay NA NA 20 128 0.01
LFL NA NA 3 128 0.01
EWC NA NA 10 256 0.001
SI NA NA 10 256 0.001
Table 7: Hyperparameter setup for Domain-Incremental setup. NA signifies Not Applicable.
Appendix H Dataset Details
H.1 Dataset summary
Datasets used for class incremental learning experiments:
  • CIFAR10, 5 tasks of 2 classes each (total of 10 classes)
  • CIFAR100 (10 tasks), 10 tasks of 10 classes each (total of 100 classes)
  • CIFAR100 (20 tasks), 20 tasks of 5 classes each (total of 100 classes)
  • TinyImageNet (40 tasks), 40 tasks of 5 classes each (total of 200 classes)
  • ImageNette (5 tasks), 5 tasks of 2 classes each (total of 10 classes)
  • ImageWoof (5 tasks), 5 tasks of 2 classes each (total of 10 classes)